OpenStudy (anonymous):

y''=\frac{ x ^{3}-2x ^{2} -x+2}{x-2}

OpenStudy (anonymous):

\[ y''=\frac{ x ^{3}-2x ^{2} -x+2}{x-2} \]

OpenStudy (anonymous):

\[y'' = ? , y=\frac{ x ^{3}-2x ^{2} -x+2}{ x-2 }\]

OpenStudy (anonymous):

divide and then integrate i guess

OpenStudy (anonymous):

oh divide and then differentiate might be easiest

OpenStudy (anonymous):

you might note that \[\frac{ x ^{3}-2x ^{2} -x+2}{ x-2 }=x^2-1\]

OpenStudy (anonymous):

then take the derivative twice, get \(2\)

OpenStudy (anonymous):

so long as \(x\neq 2\) that is

OpenStudy (anonymous):

really , thats explain the x=0 (y')